OpenStudy (iwanttogotostanford):

HELP ME HOW DO I PUT INTO MY CALCULATOR AND SOLVE THIS : What is the solution to the equation 7 ^−3x ≈ 9?

OpenStudy (tkhunny):

Have you considered graphing \(f(x) = 7^{-3x}-9\)?

OpenStudy (phi):

another way is to do some preliminary work on paper: \[ 7^{-3x}= 9 \] "Take the log (base 10 for example) " of both sides can you write down what you get ?

OpenStudy (iwanttogotostanford):

@phi I am doing a pre-worksheet for the chapter, our teacher makes us try to figure out how to do them ourselves for a grade before we actually learn the chapter,,,, so I have no idea how to do this

OpenStudy (iwanttogotostanford):

@tkhunny ^^^

OpenStudy (tkhunny):

"our teacher makes us try to figure out how to do them ourselves " Well, there you go. You had better get started.

OpenStudy (iwanttogotostanford):

@tkhunny we can use online resources goodness i know what i can and can't do

OpenStudy (phi):

"Take the log" of , for example, 3 means write log(3) if you had an equation a = b take the log of both sides means write log(a)= log(b) you do that to your equation and then, on the left side you use this property of logs: \[ \log\left( a^b\right) = b \log(a) \] to rewrite the left side.

OpenStudy (iwanttogotostanford):

ok, thank you! @phi